Solve : 2(2y^2+yz -z^2)dx+x(4y+z)dy+x(y-2z)dz= 0


1
Expert's answer
2021-07-05T14:48:03-0400

"2(2y^2+yz -z^2)dx+x(4y+z)dy+x(y-2z)dz= 0"

"2y^2+yz -z^2=(2y-z)(y+z)"

Put "2y-z=u" , "y+z=v" . Then "u+v=3y" and "2v-u=3z"

"4y+z = (4(u+v)+(2v-u))\/3 = u+2v"

"y-2z = ((u+v)-2(2v-u))\/3 = u-v"

"2(2y^2+yz -z^2)dx+x(4y+z)dy+x(y-2z)dz ="

"= 2uvdx+x(u+2v)d(u+v)\/3 + x(2v-u)d(u-v)\/3 ="

"= 2uvdx+x((u+2v) - (u-v))du\/3 + x((u+2v) +2 (u-v))dv\/3 ="

"= 2uvdx+xvdu + xudv = 0"

Special solutions: (i) "x=0" , (ii) "u=0" , (iii) "v=0"

Assuming "x\ne 0" , "u\ne 0" and "v\ne 0" , we may write then

"0 = 2uvdx+xvdu + xudv=xuv(2\frac{dx}{x} +\frac{du}{u}+\frac{dv}{v})=xuv\cdot d\log|x^2uv|"

"x^2uv=C" (the general solution)

The special solutions (i) - (iii) satisfy the equation of the general solution for C=0.


Answer. "x^2(2y-z)(y+z)=C"